# How Much are LLMs Contaminated?

### v1

@@ 3.2.1 Model Memorization
When analyzing model memorization, researchers focus on which training data points have been entirely memorized by the model, a process that can readily be leveraged to flag contamination on downstream tasks.
Elangovan et al. (2021) highlight that the common practice of random data shuffling in NLP tasks fails to account for the overlaps between training and test sets, leading to data leakage and inflated performance metrics. Models that merely memorize instead of generalize may benefit from this overlap. By employing a bag-of-words approach to assess text similarity, the study provides a foundational method for identifying and mitigating leakage, though it acknowledges the complexity of semantic similarity measurement and the potential for more sophisticated methodologies to refine overlap detection.
Lee et al. (2021) discuss the substantial benefits of deduplicating datasets for training. Deduplication significantly reduces the incidence of memorized data emission by a factor of 10×, mitigates the common issue of train-test overlap, thereby preventing overestimation of model accuracy, enhances training efficiency by reducing dataset size by up to 19%, and maintains or improves model perplexity, allowing for faster attainment of higher accuracy. This comprehensive approach not only streamlines the training process but also contributes to more environmentally and economically sustainable model development, without compromising data quality or model performance.
Magar & Schwartz (2022) find evidence of contamination across several benchmarks, impacting the performance of models like GPT-3. A novel approach to quantify the effect of contamination involves training a model on a mix of general and task-specific data, then comparing its performance on seen versus unseen instances to measure memorization and exploitation. This methodology highlights the nuanced relationship between memorization and exploitation, indicating that larger models or those trained with certain data configurations may exhibit different levels of sensitivity to contaminated data.
Gemma Team Google DeepMind (2024) emphasize the potential vulnerabilities of aligned models to adversarial attacks inducing recitation of memorized training data. Through comparable methods to Anil et al. (2023), they analyze discoverable memorization across 10k documents from the Gemma pre-trained models’ corpora. The analysis distinguishes between exact memorization—where the model’s generated text precisely matches the source—and approximates memorization, gauged by a 10% edit distance threshold. The findings reveal that Gemma models exhibit memorization rates comparable to those of the PaLM and PaLM-2 models.
3.2.2 Prompting Techniques
@@
Researchers have engineered elaborated prompting techniques that elicit data point completion by the LLM, where an output suspiciously close to the actual training data point could indicate contamination. Standard prompting asking the LLM to complete a data point may fail due to the LLM’s alignment during RLHF (Ouyang et al., 2022b).
Nasr et al. (2023) uncover the vulnerability of LLMs including ChatGPT to data extraction attacks, revealing a critical privacy concern. By prompting with an off-(training)-distribution input, they show that ChatGPT can regurgitate entire chunks of training data, including sensitive personal-level information. The study illustrates that even models trained with alignment techniques, aimed at reducing the emission of memorized data, can be induced to divulge substantial amounts of sensitive information, calling for the development of more robust defenses. The implications of this research are far-reaching, emphasizing the importance of addressing the privacy vulnerabilities inherent in the deployment of LLMs.
@@
Weller et al. (2023) highlight the potential for steering LLMs towards generating more factual and grounded content through effective prompting strategies. By appending instructional phrases such as Accoding to that encourage quoting from specific corpora, this study observes improvements in grounding as measured by the Quoted Information Precision (QUIP) Score. This method shows promise across various domains and corpora, indicating its versatility and effectiveness in leveraging LLMs’ memorized knowledge for generating more accurate and reliable responses.
3.2.3 Model Likelihood
Methods based on the likelihood utilize the fact that a model’s next token predictions are more likely to be more confident if the data point has been seen during training.
The Min-KProb technique (Shi et al., 2024) assumes white-box access to the LLM’s logits or next-token probability distributions. Given a test sample $X$, the method consists in running the LLM through all tokens of $X$, and keeping track of the $k$tokens with the smallest predicted probability. Then, the average between such bottom probabilities is computed and $X$ is deemed contaminated if this average is too high.
@@ 3.2.3 Model Likelihood
Li (2023a) compares perplexity on benchmark samples against memorized and clean baselines. The study finds significant memorization in recent models on popular reading comprehension and summarization benchmarks, while multiple-choice benchmarks show less evidence of contamination. This method provides a tool for the community to conduct rigorous contamination analysis, enabling more accurate and reliable model evaluation.
Dong et al. (2024) propose two novel likelihood-based contamination detection methodologies: CDD (Contamination Detection via output Distribution) and TED (Trustworthy Evaluation via output Distribution). CDD detects data contamination by observing the peakedness in the LLM’s output distribution in a black-box manner. It represents a significant improvement over existing approaches, offering average relative improvements of 21.8%-30.2% in terms of Accuracy, F1 Score, and AUC metrics. TED corrects the LLM’s output distribution to mitigate the impact of data contamination on evaluation metrics, significantly reducing performance improvements attributed to data contamination across various scenarios and contamination degrees.
3.3 LLM-based Model Contamination Detection
LLMs also establish a new paradigm in model contamination detection.
Golchin & Surdeanu (2023b) introduce guided prompting: unlike a standard prompt asking the LLM to complete the data point, guided prompting also includes extra information such as the dataset name. Contamination is then assessed based on the averaged difference in performance between standard and guided prompting, or if GPT-4 with in-context learning finds an exact match or two near-exact matches in the guided completions. This latter method with GPT-4 is highly accurate in identifying contamination (92%–100%). Moreover, the investigation highlights the prevalence of contamination in datasets such AG News (Zhang et al., 2015), WNLI (Levesque et al., 2012), and XSum (Narayan et al., 2018), emphasizing the critical need to address data integrity in LLM applications.
The same authors follow up with the Data Contamination Quiz (DCQ) evaluation framework for contamination detection in black-box LLMs (Golchin & Surdeanu, 2023a). The LLM under investigation is prompted with five completion options, where one option is the ground-truth text from the original dataset, three other options are paraphrases from GPT-4, and the last choice is None. The assumption is that if the LLM picks the exact answer, it is doing so out of memorization, and the authors show that this DCQ framework finds contamination cases than the guided prompting method.
@@ 4.1 Best Practices
Recent works have called the community to adopt better practices to reduce contamination.
@@
Aside from technical methodologies, Balloccu et al. (2024) conduct a systematic literature review of 225 papers and carefully detail data leakage from them to closed-source models like the GPT-series. Overall, they conclude that $\sim$42of reviewed papers leaked data to GPT-3.5 and GPT-4 for a total of $\sim$4.7M benchmark samples across 263 benchmarks. The authors also report a series of evaluation malpractices and propose a list of suggested practices for evaluating closed-source LLMs.
@@ 4.2 New Evaluation Benchmarks
Several datasets have been introduced aiming to provide contamination-free evaluation:
@@
  • WIKIMIA (Shi et al., 2024) is a dynamic benchmark made of Wikipedia events created after model training (i.e., after 2023). They further introduce the paraphrase setting where ChatGPT is leveraged to create paraphrased examples for evaluation, as also performed in CleanEval (Zhu et al., 2023).
  • KIEval (Yu et al., 2024) is an interactive evaluation framework incorporating an LLM-powered ”interactor”, which can ask follow-up questions in multi-round dialogue that leads to contamination-resilient evaluation. The evaluator could eventually determine whether a model’s response is merely a recall of benchmark answers or demonstrates a deep comprehension to apply knowledge in more complex conversations.
  • LiveCodeBench (Jain et al., 2024) continuously collects new problems from LeetCode, AtCoder and CodeForces to provide a contamination-free code generation benchmark. They show a drop in performance on new LeetCode problems for some models including DeepSeek (Bi et al., 2024), hinting at contamination.
  • Termite (Ranaldi et al., 2024) is a new text-to-SQL dataset, locked out of public access through search engines via an encryption key. Termite is made of handcrafted databases each paired with around five queries, designed to match properties of the Spider dataset (Yu et al., 2018), which unfortunately shows high signs of contamination on GPT-3.5.
Moreover, Alzahrani et al. (2024) point out that slight perturbations to existing benchmarks may disrupt model rankings on leaderboards, calling for better evaluation benchmarks.
LLM Average ARC HellaSwag MMLU TruthulQA Winogrande GSM8K
Llama-2-7B 0.278 0.064 0.489 0.146 0.001 0.001 0.968
Qwen-1.5-7B 0.432 0.000 0.023 0.700 0.664 0.202 1.000
Mistral-7B 0.312 0.000 0.000 0.010 0.042 0.999 0.444
Gemma-7B 0.417 0.978 0.000 0.580 0.122 0.820 0.000
[…]
Table 1: P-value of the guided prompting (with ROUGE-L) contamination detection (Golchin & Surdeanu, 2023b) for five LLMs on OpenLLM Leaderboard datasets. Shaded results are not statistically significant.
@@ 4.3 Future Challenges
We believe in the need for comprehensive ethical and legal frameworks (Chang, 2021) to govern the collection, usage, and management of data for LLM training. This includes policies and protocols for data privacy, consent, and use that help prevent the incorporation of contaminated data from unethical sources and the contamination of widely used pre-training data sources (e.g, CommonCrawl). Developing contamination detection techniques that do not compromise individual privacy will be crucial (Lin et al., 2010; Hayes & Ohrimenko, 2018).
@@ 5 The LLMSanitize Library
LLM Average ARC HellaSwag MMLU TruthulQA Winogrande GSM8K BBC 2024
Llama-2-7B 0.389 0.258 0.000 0.967 0.154 0.171 0.781 0.889
Qwen-1.5-7B 0.247 0.426 0.000 0.476 0.255 0.115 0.209 0.698
Mistral-7B 0.381 0.862 0.000 0.376 0.057 0.485 0.507 0.304
Gemma-7B 0.300 0.234 0.000 0.255 0.450 0.215 0.648 0.371
Table 2: P-value of the sharded likelihood test (Oren et al., 2023) for four LLMs on OpenLLM Leaderboard datasets. Shaded results are not statistically significant.
LLM Average ARC HellaSwag MMLU TruthulQA Winogrande GSM8K BBC 2024
Llama-2-7B 9.01 8.92 8.84 9.06 9.27 10.88 7.07 6.98
Qwen-1.5-7B 10.58 10.69 10.10 10.47 11.20 14.02 7.02 8.39
Mistral-7B 9.76 9.77 10.49 9.53 9.26 12.60 6.91 6.34
Gemma-7B 76.23 53.43 33.47 59.27 157.82 125.84 27.55 18.48
Table 3: Min-20% Prob (Shi et al., 2024) for four LLMs on OpenLLM Leaderboard datasets. Lower values mean greater risk of contamination.
To facilitate progress in contamination detection in LLMs, we have built and are releasing an open-source Python library, named LLMSanitize. LLMSanitize supports both data contamination and model contamination types of use cases, and input/label contamination. A minimal pseudo-code to use our library is of the following structure:
from llmsanitize import DataContaminationChecker, ModelContaminationChecker args = <…> contamination_checker = ModelContaminationChecker(args) contamination_checker.run_contamination(args.method)
The user is expected to specify the training and evaluation datasets for data contamination ; and the LLM and evaluation dataset for model contamination. Datasets and LLMs are expected to be available on HuggingFace hub, namely transformers (Wolf et al., 2020) and datasets (Lhoest et al., 2021) libraries, which our code builds upon. Efficient LLM inference for model-based use cases is handled with the vLLM library (Kwon et al., 2023). Other key parameters being controlled through arguments include the number of evaluation data points to process, and hyper-parameters specific to each contamination detection method.
We now put to test LLMSanitize. We apply three recent popular model contamination methods which we reviewed above: guided prompting (Golchin & Surdeanu, 2023b), sharded likelihood (Oren et al., 2023) and Min-KProb (Shi et al., 2024) on four widely used 7B-size LLMs: Llama-2 (Touvron et al., 2023), Qwen-1.5 (Bai et al., 2023), Mistral (Jiang et al., 2023) and Gemma (Gemma Team Google DeepMind, 2024). We use the chatbot version for each LLM, and measure contamination on the six datasets of HuggingFace OpenLLM Leaderboard: ARC (Clark et al., 2018), HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2020), TruthulQA (Lin et al., 2021), Winogrande (Sakaguchi et al., 2021) and GSM8K (Cobbe et al., 2021) ; subsampling 100 data points from each test set. We also evaluate GPT-3.5-turbo on guided prompting - the blocked access to the logprobs prohibits us from applying the other two methods. Lastly, for comparison purpose we report likelihood-based results on contamination-free data: BBC news articles from March 2024.
Results for guided prompting, sharded likelihood and Min-KProb are in Table 1, Table 2, and Table 3, respectively. The HellaSwag benchmark (Zellers et al., 2019) shows strong signs of contamination, consistent across methods and LLMs. We also report data contamination results in Appendix A and model contamination results with base models in Appendix B.

### v3

@@ 3.2 Model Memorization
An intuitive approach to detect contamination in LLMs is to frame it as a memorization problem. Researchers focus on which training data points have been memorized verbatim by the model.
Historically, this problem has been addressed in the field of privacy-preserving machine learning: adversarial attacks were crafted as membership inference attacks to access training data points memorized by a black-box language model, potentially compromising sensitive private information. We refer the reader to relevant literature in this research direction (Carlini et al., 2021, 2022).
Elangovan et al. (2021) highlight that some LLMs may benefit from an overlap between training and test sets, and merely memorize data points in this intersection, inflating their test set performance. This overlap is due to the common practice of simple random data shuffling in NLP tasks. By employing a bag-of-words approach to assess text similarity, the study provides a foundational method for identifying and mitigating leakage, though it acknowledges the complexity of semantic similarity measurement and the potential for more sophisticated methodologies to refine overlap detection. Magar and Schwartz (2022) find evidence of contamination across several benchmarks, impacting the performance of models like GPT-3. They propose to quantify the effect of contamination by training on a mix of general and task-specific data, then comparing performance on seen versus unseen instances to measure memorization and exploitation. This methodology highlights the nuanced relationship between memorization and exploitation. Gemma Team Google DeepMind (2024) emphasize the potential vulnerabilities of aligned models to adversarial attacks inducing recitation of memorized training data. Through comparable methods to Anil et al. (2023), they analyze discoverable memorization across 10k documents from the Gemma pre-trained models’ corpora. The analysis distinguishes between exact memorization — where the model’s output precisely matches the source — and approximates memorization, gauged by a 10% edit distance threshold. Gemma models exhibit memorization rates comparable to those of the PaLM and PaLM-2 models.
@@
Recently, researchers have engineered elaborated prompting techniques that elicit data point completion by the LLM, where an output suspiciously close to the actual training data point indicates memorization, which in turn means contamination. Standard prompting asking the LLM to complete a data point may fail due to the LLM’s alignment during RLHF (Ouyang et al., 2022b).
@@
By prompting with an off-(training)-distribution input, Nasr et al. (2023) show that ChatGPT can regurgitate entire chunks of training data, including sensitive personal-level information. The study illustrates that even models trained with alignment techniques, aimed at reducing the emission of memorized data, can be induced to divulge substantial amounts of sensitive information, calling for the development of more robust defenses. The implications of this research are far-reaching, emphasizing the importance of addressing the privacy vulnerabilities inherent in the deployment of LLMs. Weller et al. (2023) highlight the potential for steering LLMs towards generating more factual and grounded content through effective prompting strategies. By appending instructional phrases such as Accoding to that encourage quoting from specific corpora, this study observes improvements in grounding as measured by the Quoted Information Precision (QUIP) Score. This method shows promise across various domains and corpora, indicating its versatility and effectiveness in leveraging LLMs’ memorized knowledge for generating more accurate and reliable responses.
Cloze tasks constitute a special type of prompting where parts of the input is masked. Chang et al. (2023) design a cloze task in which the LLM is prompted with a book’s passage, from which the character name has been replaced by a [MASK] token. Correctly predicting the character’s name out of several passages per book assesses whether the LLM has seen this book during pre-training. The name cloze accuracy is very high for GPT-4 on several popular books such as Alice’s Adventures in Wonderland, and shows a striking difference with BERT’s performance on the same task. Deng et al. (2023) present the TS-Guessing method: one of the options in question-answering benchmarks test sets is hidden, and the LLM is asked to predict it. Exact match rates above 50% for OpenAI models suggest contamination on the MMLU dataset (Hendrycks et al., 2020). The work of Ranaldi et al. (2024) also uses a cloze task: they prompt the LLM to reconstruct columns names in SQL dumps, and compare the performance on an older and more recent dataset to assess contamination. GPT-3.5 accuracy falls from 33.42% to 13.21%, indicating contamination on the older Spider dataset.
3.3 Model Confidence
Going beyond simple string recitation of memorized data points, researchers may analyze the LLM’s confidence to detect contamination: a (too) high confidence may signify prior exposure to the data. Confidence analysis yields several powerful contamination detection techniques, which we highlight require white-box access to the LLM, as researchers need to obtain the output probability distribution for each token during inference.
Next-token likelihood and perplexity (by extension to a whole dataset) are widely used in contamination detection. Given a test sample $X$, the Min-KProb technique (Shi et al., 2024) consists in running the LLM through all tokens of $X$, and keeping track of the $k$tokens with the smallest predicted probability. Then, the average between such bottom probabilities is computed and $X$ is deemed contaminated if this average is too high. Li (2023a) also work on token-level probabilities and compare perplexity on benchmark samples against memorized and clean baselines. The study finds significant memorization in recent models on popular reading comprehension and summarization benchmarks, while multiple-choice benchmarks show less evidence of contamination. This method provides a tool for the community to conduct rigorous contamination analysis, enabling more accurate and reliable model evaluation. Dong et al. (2024) propose two novel likelihood-based contamination detection methodologies: CDD (Contamination Detection via output Distribution) and TED (Trustworthy Evaluation via output Distribution). CDD detects data contamination by observing the peakedness in the LLM’s output distribution in a black-box manner. It represents a significant improvement over existing approaches, offering average relative improvements of 21.8%-30.2% in terms of Accuracy, F1 Score, and AUC metrics. TED corrects the LLM’s output distribution to mitigate the impact of data contamination on evaluation metrics, significantly reducing performance improvements attributed to data contamination across various scenarios and contamination degrees. The work of Xu et al. (2024b) proposes to use perplexity and accuracy on next n-gram prediction. They compute the relative loss in performance on these metrics on paraphrased versions of the dataset created by ChatGPT compared to the original dataset. The difference of relative loss between training and test set is later used to assess contamination or benchmark leakage. Experiments on GSM8K (Cobbe et al., 2021) hint at contamination for QWen (Bai et al., 2023), Aquila and InternLM (Cai et al., 2024) series of LLMs.
@@ 3.3 Model Confidence
3.4 Model-based Closed-Data Contamination Detection
Relying on an external, high-performing LLM like GPT-4 can be useful in closed-data contamination detection, similarly to the open-data setup.
Golchin and Surdeanu (2023b) make use of GPT-4’s abilities in guided prompting, an enhanced prompting process where the completion prompt includes extra information such as the dataset name. Contamination is then assessed based on the averaged difference in performance between standard and guided prompting, or if GPT-4 with in-context learning finds an exact match or two near-exact matches in the guided completions. This latter method with GPT-4 is highly accurate in identifying contamination (92%–100% accuracy). Moreover, the investigation highlights the prevalence of contamination in datasets such as AG News (Zhang et al., 2015), WNLI (Levesque et al., 2012), and XSum (Narayan et al., 2018), emphasizing the critical need to address data integrity in LLM applications. The same authors follow up with the Data Contamination Quiz (DCQ) evaluation framework for contamination detection in black-box LLMs (Golchin and Surdeanu, 2023a). LLMs are prompted with five completion options, where one option is the ground-truth text from the original dataset, three other options are paraphrases from GPT-4, and the last choice is None. The assumption is that if the LLM picks the exact answer, it is doing so out of memorization, and the authors show that this DCQ framework finds contamination cases than the guided prompting method.
Through the use case of fine-tuning LLMs on mathematical reasoning datasets like GSM8K (Cobbe et al., 2021), Tu et al. (2024) introduce a new paradigm in (closed-data) contamination detection. Their DICE approach consists in finding the layer of the fine-tuned LLM furthest away from its counterpart layer in the not fine-tuned same LLM. Then, a MLP binary classifier is inferred on the isolated layer’s weights to predict contamination. This method is particularly effective for in-distribution contamination detection, a setup defined by the authors as fine-tuning on a dataset from the same distribution as the evaluation dataset. In this setup, DICE reaches near-perfect AUC, outperforming all other methods, including notably Min-KProb.
@@ 4.1 Best Practices
Beyond detecting contaminated evaluation data points, recent works have called the community to adopt better practices to reduce contamination.
@@
Aside from technical methodologies, Balloccu et al. (2024) conduct a systematic literature review of 225 papers and carefully detail data leakage from them to closed-source models like the GPT-series. Overall, they conclude that $\sim$42of reviewed papers leaked data to GPT-3.5 and GPT-4 for a total of $\sim$4.7M benchmark samples across 263 benchmarks. The authors report evaluation malpractices and propose a list of suggested practices for evaluating closed-source LLMs.
Several data contamination mitigation techniques are detailed in the concurrent survey work of Xu et al. (2024a).
@@ 4.2 New Evaluation Benchmarks
Several datasets have been introduced aiming to provide contamination-free evaluation.
Dynamic Evaluation
Some new datasets discard the static paradigm of existing LLM evaluation datasets, and instead propose a dynamic benchmark to avoid contamination:
@@
  • WIKIMIA (Shi et al., 2024) is a dynamic benchmark made of Wikipedia events created after model training (i.e., after 2023). ChatGPT is leveraged to paraphrase examples for evaluation, as also performed in CleanEval (Zhu et al., 2023).
  • KIEval (Yu et al., 2024) is an interactive evaluation framework incorporating an LLM-powered "interactor", which can ask follow-up questions in multi-round dialogue that leads to contamination-resilient evaluation.
  • LiveCodeBench (Jain et al., 2024) continuously collects new problems from LeetCode, AtCoder and CodeForces to provide a contamination-free code generation benchmark.
Protected Evaluation
Another strategy to avoid contamination of evaluation data is to protect it by locking it out of the public domain. Termite (Ranaldi et al., 2024) is a new text-to-SQL dataset, locked out of public access through search engines via an encryption key. Termite is made of handcrafted databases, each paired with around five queries, designed to match properties of the Spider dataset (Yu et al., 2018), which shows high signs of contamination on GPT-3.5.
Alternative Evaluation
New approaches to LLM evaluation are emerging in the era of widespread public benchmark contamination. The DCQ benchmark utilizes GPT-4 to paraphrase data points from famous benchmarks such as WNLI (Levesque et al., 2012), and checks whether the LLM picks the exact continuation or a paraphrased version, where the former strongly suggests contamination. Moreover, Alzahrani et al. (2024) point out that slight perturbations to existing benchmarks may disrupt model rankings on leaderboards, calling for better evaluation benchmarks.
@@ 4.3 Future Challenges
We believe in the need for comprehensive ethical and legal frameworks (Chang, 2021) to govern the collection, usage, and management of data for LLM training. This includes policies and protocols for data privacy, consent, and use that help prevent the incorporation of contaminated data from unethical sources and the contamination of widely used pre-training data sources (e.g., CommonCrawl). Developing contamination detection techniques that do not compromise individual privacy will be crucial (Lin et al., 2010; Hayes and Ohrimenko, 2018).
@@ 5 The LLMSanitize Library
LLM ARC HellaSwag MMLU TruthulQA Winogrande GSM8K
Guided Prompting (ROUGE-L)
Llama-2-7B 0.064 0.489 0.146 0.001 0.001 0.968
Qwen-1.5-7B 0.000 0.023 0.700 0.664 0.202 1.000
Mistral-7B 0.000 0.000 0.010 0.042 0.999 0.444
Gemma-7B 0.978 0.000 0.580 0.122 0.820 0.000
GPT-3.5-turbo 0.000 0.184 0.970 0.075 0.000 1.000
Sharded Likelihood
Llama-2-7B 0.258 0.000 0.967 0.154 0.171 0.781
Qwen-1.5-7B 0.426 0.000 0.476 0.255 0.115 0.209
Mistral-7B 0.862 0.000 0.376 0.057 0.485 0.507
Gemma-7B 0.234 0.000 0.255 0.450 0.215 0.648
Table 1: P-value for two closed-data contamination detection methods using a statistical test: Guided Prompting (Golchin and Surdeanu, 2023a) (top) and Sharded Likelihood (Oren et al., 2023) (bottom). Shaded results are not statistically significant. We ran LLMSanitize on 100 test data points.
To facilitate progress in contamination detection in LLMs, we have built and are releasing an open-source Python library, named LLMSanitize. LLMSanitize supports most of the open-data and closed-data contamination detection described above and listed in Figure 1. It relies on the HuggingFace hub to access LLMs and the data, namely transformers (Wolf et al., 2020) and datasets (Lhoest et al., 2021) libraries. Efficient LLM inference for model-based use cases is handled with the vLLM library (Kwon et al., 2023). LLMSanitize is straightforward to install through PyPI. Users are expected to pass in the relevant arguments for each contamination detection method, as well as key parameters such as the LLM and dataset name and the number of data points to run the contamination detection on. We refer to Appendix A.
We now put to test LLMSanitize. We apply Guided Prompting (Golchin and Surdeanu, 2023b) and Sharded Likelihood (Oren et al., 2023) on the six datasets of HuggingFace OpenLLM Leaderboard [^1] https://huggingface.co/open-llm-leaderboard. We employ four widely used 7B-size LLMs: Llama-2 (Touvron et al., 2023), Qwen-1.5 (Bai et al., 2023), Mistral (Jiang et al., 2023) and Gemma (Gemma Team Google DeepMind, 2024). We also evaluate GPT-3.5-turbo on guided prompting - the blocked access to the logprobs prohibits us from applying the other method. Results are shown in Table 1. The HellaSwag benchmark (Zellers et al., 2019) shows strong signs of contamination, consistent across methods and LLMs.